# Coalescence in Markov chains

### v1

@@ 1. Introduction
We close with a remark. The general questions approached here seem quite fundamental to a complete theory of finite state-space Markov chains, and yet they have received only limited attention so far. This is reflected in the partial state of knowledge revealed in [8] and the current paper. This work may be viewed as a contribution to the theory of iterated random functions.
@@ 2. Grand couplings of Markov chains
We will make frequent use of the following well known fact due to Doeblin [5] (see also [9, p. 260]).
@@ 3. Coalescence of trajectories
The relation $\sim$ is a (random) equivalence relation.
@@
It suffices to prove the transitivity of $\sim$. Let $i\sim j$ and $j\sim k$. Since $S$ is finite, there exists $T$ such that $X_{t}^{i}=X_{t}^{j}=X_{t}^{k}$ for $t\geq T$. The claim follows ∎
The equivalence classes of $\sim$ are termed the coalescence classes of $(X^{i}:i\in S)$. The number of coalescence classes is a.s. constant (see [8, Lemma 2]), and we denote this number by $k(\mu)$ and call it the coalescence number of $\mu$. We define
@@
and backward coalescence is said to occur if $\mathbb{P}(C<\infty)$. On the event $\{C<\infty\}$, ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ may be regarded as a random state, and the main theorem of CFTP asserts that, if backward coalescence occurs, then ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ is distributed as the invariant distribution of the transition matrix $P$.
@@
Recall that, since $P$ is finite and irreducible, it has a unique invariant distribution (see, for example, [9, Thm 6.4.3]).
@@
    Let $m\in\mathbb{N}$, and suppose there exists $s\in S$ such that $p_{i,s}>1/m$ for all $i\in S$, then $k_{\text{\rm max}}<m$.
@@
Whereas $k(\mu)$ is a.s. constant, the coalescence classes of $\sim$ need not themselves be a.s. constant. Here is an example of this, preceded by some notation.
@@
where $\bm{\alpha}$} is a probability mass function on $\mathcal{F}_{S}$ with support $\text{\rm supp}(\mu)$, and $\delta_{f}$ is the Dirac delta-mass on the point $f\in\mathcal{F}_{S}$. Thus, $\alpha_{f}>0$ if and only if $f\in\text{\rm supp}(\mu)$. We denote by $M(\mathcal{F}_{S})$ the set of all probability measures $\mu$ of the form (3.7).
In advance of stating an extension of [8, Lemma 3], we remind the reader of the definition of a cyclic class.
@@
(a) This is found at the proof of [8, Lemma 3], but is included here for clarity. The independence coupling of Example 2.2 gives rise to $n=|S|$ chains with transition matrix $P$, starting from $1,2,\dots,n$, respectively, that evolve independently until they meet. If $P$ is aperiodic (and, by assumption, irreducible) then all $n$ chains meet a.s. in finite time. The last holds by Theorem 2.5, since any paths that meet will remain together thenceforth under this coupling.
@@ 4. Lumpability and block measures
There is a limited literature on lumpable chains, for which the reader may consult, for example, [6, 7, 12] and [9, Exer. 6.1.13]. Kemeny and Snell [11] proved a necessary and sufficient condition for $X$ to be $\mathcal{S}$-lumpable, namely the following.
@@ A Markov chain $X$ with transition matrix $P$ is $\mathcal{S}$-lumpable if and only if,
Recall the set $M(\mathcal{F}_{S})$ of probability measures of the form (3.7).
@@
    If $\mu\in\mathcal{L}_{P}$ is an $\mathcal{S}$-block measure, then, for $r,s\in I$,
    (4.6) $\text{$\lambda_{r,s}:=\lambda_{r,s}^{(i)}$ is constant for $i\in S_{r}$}.$
@@
    The ‘block-transition matrix’ $\Lambda=(\lambda_{r,s}:1\leq r,s\leq\ell)$ is irreducible. Moreover, $\Lambda$ is doubly stochastic, which may be expressed as
@@ 5. Existence of block measures
Let $\mu$ be a probability measure with $\text{\rm supp}(\mu)=\widehat{\mathcal{F}}$. Then $p_{i,j}:=\sum_{f:f(i)=j}\mu(\{f\})>0$ for every pair $i$, $j$, whence $P=(p_{i,j})$ is irreducible. The claim follows by assumptions (i) and (ii), and Definition 4.2. ∎
@@
Let the pair $P\in\mathcal{P}_{S}$ and $\mathcal{S}=\{S_{r}:r\in I\}$ (a partition of the state space $S$ with $I=\{1,2,\dots,\ell\}$) satisfy (4.6) and be such that
(5.1) $\Lambda$ is doubly stochastic.
By Theorem 4.3 these two conditions are necessary for the existence of an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$.
@@
Let $P\in\mathcal{P}_{S}$, and let $\mathcal{S}=\{S_{r}:r\in I\}$ be a partition of $S$ with index set $I=\{1,2,\dots,\ell\}$. Assume (4.6) and (5.1) hold. The $(P,\mathcal{S},\rho)$-product measure $\mu$ satisfies $k(\mu)=\ell$ if and only if [^2] One may work with any given value of $r$ in (5.5), and we have chosen $r=1$ for concreteness. Condition (5.5) is similar to the sufficient condition of [17, Thm 6] for the quenched ergodicity of a Markov chain with random transition matrices. Theorem 5.3 corrects an error in [8, Thm 6].
@@ 7. Functions that generate transition matrices
In this final section we pose an inverse question. As usual, the state space is $S=\{1,2,\dots,n\}$. Recall the set $\mathcal{P}_{S}$ of transition matrices on $S$ and the set $\mathcal{F}_{S}$ of functions from $S$ to $S$. Given a subset $\mathcal{G}\subseteq\mathcal{F}_{S}$, let
@@
Some very special cases admit precise answers to Question 7.1. Let $\mathcal{D}_{S}\subset\mathcal{P}_{S}$ denote the set of doubly stochastic matrices, and let $\mathcal{F}^{\text{\rm perm}}_{S}\subset\mathcal{F}_{S}$ denote the set of permutations of $S$.
@@
Secondly, in Example 7.2, we have $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$ since $\mathcal{G}_{S}$ is the set of stochastic matrices with off-diagonal elements summing to $1$.
@@
It is elementary that if $\mathcal{G}$ is missing two of them then $\mathbb{Q}(\mathcal{G})=0$. The following lemma shows that (when $n=2$) removing a single function one retains positive measure.
@@ Acknowledgements
The authors are grateful to Tabet Aoul Alaa Eddine Islem for indicating an error in an earlier version of Theorem 4.3 as published in [8].

### v2

@@ 1. Introduction
The general questions approached here seem quite fundamental to a complete theory of finite state-space Markov chains, and yet they have received only limited attention so far. This is reflected in the partial state of knowledge revealed in [10] and the current paper, which concentrate on grand couplings based on sequences of independent random functions. Thus, this work may be viewed as a contribution to the theory of iterated random functions. The idea of studying a ‘forward’ chain by running it ‘backwards’ appeared in [14]; see also [5, 6], for example.
There are a number of questions and open problems presented in this paper, including Questions 3.3, 3.8, 3.11, 7.4, and the authors hope that readers will resolve some of these.
@@ 2. Grand couplings of Markov chains
We will make frequent use of the following well known fact due to Doeblin [7] (see also [11, p. 260]).
@@ 3. Coalescence of trajectories
The relation $\sim$ is a (random) equivalence relation on $S$.
@@
It suffices to prove the transitivity of $\sim$. Let $i\sim j$ and $j\sim k$. For $t\geq\max\{T_{i,j},T_{j,k}\}$ we have that $X_{t}^{i}=X_{t}^{j}=X_{t}^{k}$. The claim follows. ∎
The equivalence classes of $\sim$ are termed the coalescence classes of $(X^{i}:i\in S)$. The number of coalescence classes is a.s. constant (see [10, Lemma 2]), and we denote this number by $k(\mu)$ and call it the coalescence number of $\mu$. We define
@@
and backward coalescence is said to occur if $\mathbb{P}(C<\infty)=1$. On the event $\{C<\infty\}$, ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ may be regarded as a random state, and the main theorem of CFTP asserts that, if backward coalescence occurs, then ${\overset{{}_{\shortleftarrow}}{F}}_{C}$ is distributed as the invariant distribution of the transition matrix $P$.
@@
Recall that, since $P$ is finite and irreducible, it has a unique invariant distribution (see, for example, [11, Thm 6.4.3]).
@@
    Let $m\in\mathbb{N}$, and suppose there exists $s\in S$ such that $p_{i,s}>1/m$ for all $i\in S$. Then $k_{\text{\rm max}}<m$.
@@
By Lemma 3.13(a) below, we have that $1\in K(P)$ for $\mathbb{Q}$-a.e. $P$.
Although $k(\mu)$ is a.s. constant, the coalescence classes of $\sim$ need not themselves be a.s. constant. Here is an example of this, preceded by some notation.
@@
where $\bm{\alpha}$} is a probability mass function on $\mathcal{F}_{S}$ with support $\text{\rm supp}(\mu)$, and $\delta_{f}$ is the Dirac delta-mass on the point $f\in\mathcal{F}_{S}$. Thus, $\alpha_{f}>0$ if and only if $f\in\text{\rm supp}(\mu)$. We denote by $M(\mathcal{F}_{S})$ the set of all probability measures on $\mathcal{F}_{S}$, noting that they may be represented in the form (3.7).
In advance of stating an extension of [10, Lemma 3], we remind the reader of the definition of a cyclic class.
@@
(a) This is found at the proof of [10, Lemma 3], but is included here for clarity. The independence coupling of Example 2.2 gives rise to $n=|S|$ chains with transition matrix $P$, starting from $1,2,\dots,n$, respectively, that evolve independently until they meet. If $P$ is aperiodic (and, by assumption, irreducible) then all $n$ chains meet a.s. in finite time. The last holds by Theorem 2.5, since any paths that meet will remain together thenceforth under this coupling.
@@ 4. Lumpability and block measures
There is a limited literature on lumpable chains, for which the reader may consult, for example, [8, 9, 15] and [11, Exer. 6.1.13]. Kemeny and Snell [13] proved a necessary and sufficient condition for $X$ to be $\mathcal{S}$-lumpable, namely the following.
@@ A Markov chain $X$ with transition matrix $P$ is $\mathcal{S}$-lumpable if and only if,
Recall from (3.7) the set $M(\mathcal{F}_{S})$ of probability measures on $\mathcal{F}_{S}$.
@@
If there exists an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$, then
@@
    for $r,s\in I$,
    (4.6) $\text{$\lambda_{r,s}:=\lambda_{r,s}^{(i)}$ is constant for $i\in S_{r}$},$
@@
    the ‘block-transition matrix’ $\Lambda=(\lambda_{r,s}:1\leq r,s\leq\ell)$ is irreducible. Moreover,
    (4.7) $\Lambda$ is doubly stochastic,
    which may be expressed as
@@ 5. Existence of block measures
Let $\mu$ be a probability measure with $\text{\rm supp}(\mu)=\widehat{\mathcal{F}}$. By assumption (i), for all $r\in I$ and $j\in S_{r}, ii$ and $j$ coalesce in one step. The claim follows by assumption (ii) and Definition 4.2. ∎
@@
Let the pair $P\in\mathcal{P}_{S}$ and $\mathcal{S}=\{S_{r}:r\in I\}$ (a partition of the state space $S$ with $I=\{1,2,\dots,\ell\}$) satisfy (4.6) and (4.7). By Theorem 4.3 these two conditions are necessary for the existence of an $\mathcal{S}$-block measure $\mu\in\mathcal{L}_{P}$.
@@
Let $P\in\mathcal{P}_{S}$, and let $\mathcal{S}=\{S_{r}:r\in I\}$ be a partition of $S$ with index set $I=\{1,2,\dots,\ell\}$. Assume (4.6) and (4.7) hold. The $(P,\mathcal{S},\rho)$-product measure $\mu$ satisfies $k(\mu)=\ell$ if and only if [^2] One may work with any given value of $r$ in (5.4), and we have chosen $r=1$ for concreteness. Condition (5.4) is similar to the sufficient condition of [20, Thm 6] for the quenched ergodicity of a Markov chain with random transition matrices. Theorem 5.3 corrects an error in [10, Thm 6].
@@ 7. Functions that generate transition matrices
In this final section we pose an inverse question. As usual, the state space is $S=\{1,2,\dots,n\}$. Recall the set $\mathcal{P}_{S}$ of irreducible transition matrices on $S$ and the set $\mathcal{F}_{S}$ of functions from $S$ to $S$. Given a subset $\mathcal{G}\subseteq\mathcal{F}_{S}$, let
@@
Some further very special cases admit precise answers to Question 7.1. Let $\mathcal{D}_{S}\subset\mathcal{P}_{S}$ denote the set of doubly stochastic matrices, and let $\mathcal{F}^{\text{\rm perm}}_{S}\subset\mathcal{F}_{S}$ denote the set of permutations of $S$.
@@
Secondly, in Example 7.2, we have $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$ since $\mathcal{G}$ is the set of stochastic matrices with off-diagonal elements summing to $1$.
@@
It is elementary that if $\mathcal{G}$ is missing two of them then $\mathbb{Q}(\mathcal{P}(\mathcal{G}))=0$. The following lemma shows that (when $n=2$) removing a single function one retains positive measure.
@@ Acknowledgements
The authors are grateful to Tabet Aoul Alaa Eddine Islem for indicating an error in an earlier version of Theorem 4.3 as published in [10]. They thank two referees for their useful comments.